A revolver of mass 5kg releases a bullet of mass 0.10kg at a speed of 100m/s.find the revolvers speed recoil
1
Expert's answer
2020-01-20T07:02:53-0500

Mass of revolver "M=5kg"

Mass of bullet "m=0.10kg"

Velocity of bullet "v=100m\/s"

Velocity of revolver "V-?"


Solution:

"v_0=0\;m/s" velocity of revolver and bullet (before)

Momentum of the system (before) "p_0=(M+m)v_0=0"

Momentum of the system (after) "p=MV+mv"

"p_0=p"

"0=MV+mv"

"0=5V+0.10*100"

"V=-2\; m/s"

The negative sign indicates the revolver recoil backwards.